Duhwa is 4 years more than 22 times as old Amy. If the sum of their ages is 96, how old is Duhwa?
Answer by checkley77(12844)

D=22A+4
D+A=96
22A+4+A=96
23A=96-4
23A=92
A=92/23
4 YEARS OLD IS AMY'S AGE
DUHWA'S AGE=22*4+4=88+4=92 YEARS OLD IS DUHWA'S AGE.
PROOF:92+4=96
96=96